a 210 yard long rope is cut into 3 pieces. the first piece is 3 times as long as the 2nd piece. the third piece is 2 times as long as the 2nd piece. what is the length of the longest piece of rope?
Answer by mananth(16949)

let the second piece be x
the first piece is 3 times second. = 3x
the third piese is twice the second = 2x
total length = 210 yards.
..
the sum of lengths is 210
x+2x+3x=210
6x=210
x=210/6
x=35
the longest piece = 3x = 3*35=105 yards